Help please
Assignment Assignment> open Assignment ASSIGNMENT R 3.4 Problem 3.059 dPractice Reading Incorrect An Dom The following function is probability mass function. f (x) teview Score Determine the mean, μ, and variance, σ, of the random variable. Enter the exact answers (as fractions if necessary) Oblective Moble Site Statistical Tables and Charts Click if you would like to Show Work fer this questien Qn Shom Weck Question Attempts: 1 of 3 used SAVE FOR LATER SUBMIT ANSWER
Assignment Assignment> open Assignment ASSIGNMENT R 3.4 Problem 3.060 dPractice Reading Determine the mean and variance of the random variable with the following probability mass function. f (x) (64/21)4)x-12,3 Round your answers to three decimal places (e.g. 98.765) Mean = teview Score Oblective Moble Site Click if you would like to Show Work for this question: Oen Shom Wack question Attempts: 0 of 3 used SAVE FOR LATER SUBMIT ANSWER 4 Maple,et
Answer #1

3.059)

x P(x) xP(x) x2P(x)
-2 1/8 -0.250 0.500
-1 1/4 -0.250 0.250
0 1/4 0.000 0.000
1 1/4 0.250 0.250
2 1/8 0.250 0.500
total 0.000 1.500
E(x) =μ= ΣxP(x) = 0.0000
E(x2) = Σx2P(x) = 1.5000
Var(x)=σ2 = E(x2)-(E(x))2= 1.500

from above mean μ=0

and variance σ2 =3/2

3.060)

x P(x) xP(x) x2P(x)
1 16/21 0.762 0.762
2    4/21 0.381 0.762
3    1/21 0.143 0.429
total 1.286 1.952
E(x) =μ= ΣxP(x) = 1.2857
E(x2) = Σx2P(x) = 1.9524
Var(x)=σ2 = E(x2)-(E(x))2= 0.299

mean μ=1.286

variance σ2 =0.299